OpenStudy (anonymous):

WILL MEDAL & FAN !!! PLEASE HELP ASAP !!! Calculate the area of a trapezoid. The measure of the first base is 15 feet, the measure of the second is 8 feet, while the height is 4 feet. a. 26 square feet b. 36 square feet c. 46 square feet d. 56 square feet

OpenStudy (anonymous):

A=a+b/2(h)

OpenStudy (anonymous):

15+8=23

OpenStudy (anonymous):

23/2=11.5

OpenStudy (anonymous):

11.5*4=

OpenStudy (anonymous):

What does 11.5 times 4 equal?

OpenStudy (anonymous):

C . 46

OpenStudy (anonymous):

Yep! Nice job